What is the domain of f/g? if f(x) = x^2-4x-5 and g(x)=x-5

What is h *g when h = x+1/x-1 and g is x-5
1
Expert's answer
2014-10-29T13:51:21-0400
f/g = (x^2-4x-5)/(x-5) = (x+1)(x-5)/(x-5) = x+1, x is not equal to 5So, the domain of f/g is (-∞, 5) ∪ (5, +∞) (all numbers except 5)h*g = (x+1/x-1)(x-5) = ((x+1)(x-5))/(x-1) = (x^2-4x-5)/(x-1) = x - 3 - 8/(x-1)